OpenStudy (anonymous):

What is the pattern rule for 1,2,4,8,16

OpenStudy (anonymous):

Are you sure of the last number? I cannot imagine any pattern rule that includes 17 in this series.

OpenStudy (anonymous):

Can you get from 1 to 2 using addition? Will adding that same number to 2 get you to 4? Use the same thinking process with multiplication. You can eliminate subtraction and division from consideration because the numbers are increasing.

OpenStudy (anonymous):

FIRST=1=2^0=2^1-1 SECOND=2^(2-1) THIRD=2^(3-1) ........................ NTH=2^(n-1)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."